OpenStudy (anonymous):

write the equation of the line that passes through 2,3 and is parallel to the line 12x-5y=2

OpenStudy (anonymous):

i can give u the anwers they gave me if u want

OpenStudy (ankit042):

Y=mx+c where m is the slope. Now find the slope from first equation. Next to find the new equation y-y1=m(x-x1)

OpenStudy (ankit042):

@bjones5454 you need more hints?

OpenStudy (anonymous):

these are the anwers a y=12/5x-9/5 b y=5/12x-9/5 cy= 5/12x+9/5 d y=-5/12x-9/5

OpenStudy (anonymous):

@ankit042

OpenStudy (ankit042):

the value of m =12/5 y-3 = 12/5(x-2) solve to get the answer

OpenStudy (anonymous):

is it A ????

OpenStudy (anonymous):

thank u

OpenStudy (ankit042):

yes it is A ...you are welcom!